On July 1, 2020, Metlock Inc. made two sales:

1. It sold excess land in exchange for a four-year, non–interest-bearing promissory note in the face amount of $1,244,650. The land’s carrying value is $630,000.
2. It rendered services in exchange for an eight-year promissory note having a face value of $410,000. Interest at a rate of 2% is payable annually.


The customers in the above transactions have credit ratings that require them to borrow money at 12% interest. Metlock recently had to pay 6% interest for money it borrowed from British Bank.

3. On July 1, 2020, Metlock also agreed to accept an instalment note from one of its customers in partial settlement of accounts receivable that were overdue. The note calls for four equal payments of $20,700, including the principal and interest due, on the anniversary of the note. The implied interest rate on this note is 8%.


The tables in this problem are to be used as a reference for this problem.

Prepare the journal entries to record the three notes receivable transactions of Metlock Inc. on July 1, 2020. (For calculation purposes, use 5 decimal places as displayed in the factor table provided. Round answers to 0 decimal places, e.g. 58,971. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

1.

2.

3.

Prepare an instalment note receivable schedule for the instalment note obtained in partial collection of accounts receivable. (Round answers to 0 decimal places, e.g. 58,971.)

Instalment Note Receivable Schedule
Date Cash
Collected
Interest
Revenue
Principal
Collected
Note Carrying
Amount
July 1 2020 $
July 1 2021 $ $ $
July 1 2022
July 1 2023
July 1 2024
Answer #1
No Date Account titles and explanation Debit Credit
1 July 1,2020 Note receivable 1244650
Land 630000
Gain on sale of land (1244650-630000) 614650
(Land sold in exchange of note)
2 July 1,2020 Note receivable (Note:1) 257238
Sales revenue 257238
(Sales revenue recorded)
3 July 1,2020 Note receivable (Note:2) 68560
Accounts receivable 68560
(Installment note accepted for partial settlement)
Note:1
Face value of note=$ 410000
It is received at the time of maturity
Now,let's find the present value of face value of bonds
Discount factor=Interest on money borrowed=6%
Present value of face value of notes=Face value of the bonds*Discount factor at 6% for the 8th year=410000*0.62741=$ 257238
Note:2
Equal payment=$ 20700
Discount factor=Implied interest rate=8%
Present value of note=Equal payment*Discount factor at 8% for 4 years=20700*3.3121=$ 68560
Installment note receivable schedule
Date Cash collected Interest revenue Principal collected Note carrying amount
July 1,2020 68560
July 1,2021 20700 5485 15215 53345
July 1,2022 20700 4268 16432 36912
July 1,2023 20700 2953 17747 19165
July 1,2024 20700 1535 19165 0
Interest revenue=Beginning note carrying amount*Implied interest rate
Principal collected=Cash collected-Interest revenue
Ending note carrying amount=Beginning note carrying amount-Principal collected
I appreciate your ratings